# A Comprehensive Survey of Direct Preference Optimization: Datasets, Theories, Variants, and Applications

### v2

@@ 2.1 Reinforcement Learning from Human Feedback
The typical RLHF pipeline for large language models (LLMs) generally consists of three key phases: (1) supervised fine-tuning (SFT), (2) preference sampling and reward model training, and (3) reinforcement learning (RL) optimization.
In the SFT stage, RLHF typically starts with a pre-trained language model, which is further refined using supervised fine-tuning on a curated dataset of high-quality human-generated responses. This supervised fine-tuning stage produces an initial model, denoted as $\pi^{SFT}$, which has improved but still not fully optimized alignment with human preferences. This model serves as the baseline for the subsequent RLHF stages. The preference learning stage involves collecting human feedback in the form of preference data. Given pairs of responses generated by $\pi^{SFT}$, human evaluators express a preference for one response over the other. These preferences serve as the foundational feedback for reward model training. The reward model’s objective is to quantify these preferences by assigning a numerical score to each output, effectively converting human feedback into a scalar reward.
One widely used formula in this stage is the Bradley-Terry (Bradley and Terry, 1952b) (BT), which proposes a mechanism for modeling pairwise comparisons and assigning concrete reward values to the outputs. In the context of RLHF, BT is used to predict human preferences between pairs of responses. The resulting reward function is then used to guide the concrete model’s decision-making by providing feedback on how well each generated response aligns with human preferences.The BT formula proposes that the human preference distribution $p^{*}$ for a given response pair can be expressed as the following formula:
@@ 2.2 Direct Preference Optimization
In RLHF, the process is intricate, involving the training of a reward model and the iterative sampling from the language model’s policy during the training loop. This complexity arises from the need to continuously evaluate and refine the model based on feedback from the reward model, resulting in significant computational demands. Direct Preference Optimization (DPO) offers a streamlined alternative by optimizing the same objective as RLHF but bypasses the explicit need for a separate reward model, thereby reduceing the computational costs associated with aligning the LLM.
@@
This summation ensures that $\pi_{r}(y\mid x)$ is a valid probability distribution, but it is unpratical to enumerate all possible responses to traverse of all possible input x, which requires significant computational resources, especially when handling a large response space. Actually, DPO algorithm mitigates the need for explicit reward model training and sampling from the LM policy. By directly optimizing preferences without the intermediate step of training a reward model, DPO simplifies and reduces the overall computational burden, making it a more efficient approach for language model alignment. To be specific, the objective function in (4) can be rearranged to isolate $r(x,y)$, yielding:
@@
The DPO loss function is derived directly from the reward differences and avoids the need for a separate reward model. By using this approach, the policy $\pi_{\theta}$ can be optimized analytically based on the reward function differences, simplifying the optimization process and eliminating the need for an intermediate reward model. This makes DPO computationally more efficient, as it directly leverages preference data to guide the model’s learning without the added complexity of reward model training.
@@ 3 Research Questions and Variants
PPO (Ouyang et al., 2022a) $r_{\phi}(x,y)-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref }}(y\mid x)\right]$
DPO (Rafailov et al., 2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
IPO (Azar et al., 2023) $\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2}\right)^{2}$
@@
RRHF (Yuan et al., 2023) $\sum_{r_{i}<r_{j}}\max(0,p_{i}-p_{j})-\sum_{t}\log P_{\pi}(y_{i^{\prime},t}|x,y_{i^{\prime},<t})$
@@
SPIN (Chen et al., 2024a) $\ell\left(\lambda\log\frac{p_{\theta}(y|x)}{p_{\theta_{t}}(y|x)}-\lambda\log\frac{p_{\theta}(y^{\prime}|x)}{p_{\theta_{t}}(y^{\prime}|x)}\right)$
@@
Step-DPO (Lai et al., 2024) $-\log\sigma(\beta\log\frac{\pi_{\theta}(s_{w}|x,s_{1\sim k-1})}{\pi_{ref}(s_{w}|x,s_{1\sim k-1})}-\beta\log\frac{\pi_{\theta}(s_{l}|x,s_{1\sim k-1})}{\pi_{ref}(s_{l}|x,s_{1\sim k-1})})$
T-DPO (Zeng et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
KTO (Ethayarajh et al., 2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)-\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$
@@
$\beta$-DPO (Wu et al., 2024a) $-\log\sigma\left(\beta_{i}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{i}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
CPO (Xu et al., 2024b) $-\log p_{\theta}(y_{w}|x)-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)$
ORPO (Hong et al., 2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$
@@
OPtune (Chen et al., 2024b) $-R(x,y_{w},y_{l})\cdot\log\sigma\left(\beta_{1}\log\frac{\pi_{t+1}(y_{w}|x)}{\pi_{t}(y_{w}|x)}-\beta_{1}\log\frac{\pi_{t+1}(y_{l}|x)}{\pi_{t}(y_{l}|x)}\right)$
@@
IRPO  (Pang et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(c_{w},y_{w}|x_{i})}{\pi_{t}(c_{w},y_{w}|x_{i})}-\beta\log\frac{\pi_{\theta}(c_{l},y_{l}|x_{i})}{\pi_{t}(c_{l},y_{l}|x_{i})}\right)-\alpha\frac{\log\pi_{\theta}(c_{w},y_{w}|x_{i})}{|c_{w}|+|y_{w}|}$
@@
R-DPO (Park et al., 2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
LD-DPO (Liu et al., 2024d) $-\log\sigma\left(\beta_{i}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{i}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
SimPO (Meng et al., 2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$
@@
SPO (Lou et al., 2024a) $-\log\sigma\left(\xi_{2}\phi_{2}(x,y_{w},y_{l})-\xi_{1}\phi_{1}(x,y_{w},y_{l})\right)$
@@
Table 1: Part variants of DPO objectives discussed in RQs. Note that a method may be discussed across multiple RQs. The table is inspired from  (Meng et al., 2024).
@@
RQ0: why DPO? Large Language Models have recently developed rapidly with the supervised fine-tuning (SFT) training (Chung et al., 2022). However, due to the next token prediction misaligned with the objective "follow the user’s instructions helpfully and safely" (Brown et al., 2020; Fedus et al., 2022), SFT models often exhibit unintended behaviors. For example, they might generate toxic, harmful, hallucinatory, and biased responses (Chung et al., 2022). Furthermore, SFT solely instructs models to learn from correct responses while neglecting to address incorrect ones, thus lacking the consideration of human values or preferences (Zhao et al., 2023).
@@
A potential solution to this issue is Reinforcement Learning from Human Feedback (RLHF), using human preferences as a reward signal to fine-tune models by reinforcement learning (Chung et al., 2022; Bai et al., 2022b; Lee et al., 2023), such as PPO (Schulman et al., 2017b) and actor-critic (Mnih et al., 2016; Haarnoja et al., 2018). We follow the mainstream work (Ziegler et al., 2019; Stiennon et al., 2020) and focus on PPO (Schulman et al., 2017b) in this paper.
@@
Nevertheless, RLHF requires loading four models (policy, reference, reward, and critic) and involves training sensitive hyperparameters along with substantial and costly human oversight to develop a well-trained reward model. Furthermore, some research has identified several challenges associated with reward modeling. For instance, reward misspecification  (Pan et al., 2022) occurs when a scalar score from the reward model fails to comprehensively represent human preferences. Additionally, misgeneralization  (Tien et al., 2023) arises when reward models compute rewards using unexpected or contingent features of the environment  (Michaud et al., 2020), leading to causal confusion and poor out-of-distribution generalization. Another issue is reward hacking  (Casper et al., 2023), for example, without regularization to penalize the KL divergence between a base model and the fine-tuned model, large language models (LLMs) undergoing RL often learn to output nonsensical text  (Stiennon et al., 2020).
@@
A powerful alternative that does not require an explicit reward model or RL to RLHF is Direct Preference Optimization (DPO) (Rafailov et al., 2023). DPO derived the closed-form solution of the PPO optimization objective, revealing the relationship between the reward and the optimal policy model. It then reparameterizes the reward modeling loss with the policy, thereby calculating implicit rewards and optimizing the policy model $\pi_{\theta}$ over the preference data as depicted in EQ.8.
@@
Compared to RLHF, DPO is empirically stable, performant, and computationally lightweight (Rafailov et al., 2023; Ethayarajh et al., 2024). However, recent studies indicate that, despite avoiding explicit reward modeling, DPO faces challenges similar to those encountered in RLHF. For example, the offline preferences can lead DPO to overfit the training data distribution, thereby limiting its generalization capabilities (Azar et al., 2023). DPO is sensitive to the distribution shift between the base model outputs and preference data. Besides, DPO is prone to generating a biased policy that favors out-of-distribution responses, leading to unpredictable behaviors, etc.  (Xu et al., 2024a; Saeidi et al., 2024).
@@
RQ1: Effect of Implicit Reward Modeling. We first discuss the generalization ability of the implicit reward modeling in DPO. DPO avoids training an explicit reward model by establishing a mapping from reward functions to optimal policies. This enables us to transform a loss function over reward functions into a loss function over policies. $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ is the reward implicitly defined by the policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$.
@@
IPO (Azar et al., 2023) points out the overfitting reward problem in DPO and proposes a novel loss to this issue.  Li et al. (2024a) conducted a theoretic and empirical analysis of the errors inherent in policy optimization methods when learning from user preferences for alignment in reinforcement learning including PPO, DPO, and IPO. Their experimental results underscore the critical importance of optimizing policies using out-of-distribution preference data. Furthermore, the study demonstrates the effectiveness of employing an explicit reward model to enhance policy performance.
@@
Dong et al. (2023) proposed a reward ranked fine-tuning method to explore the list-wise feedback. The core idea of RAFT is that the model iteratively learns from the induced best-of-K policy  (Nakano et al., 2022; Cobbe et al., 2021), which samples K responses and selects the one with the highest reward as the final output. Then the model is fine-tuned on the optimal responses.
Rank Responses to align Human Feedback (RRHF)  (Yuan et al., 2023) fully exploited rank from human annotators or reward models by combining a modified rank loss with SFT loss. To avoid explicit reward model, they take length-normalized conditional log probability of responses under policy model $\pi_{\theta}$ as reward score. The core idea is to let the policy model $\pi_{\theta}$ give larger probabilities for better responses and give smaller probabilities for worse responses.
@@
Besides,  Ethayarajh et al. (2024) proposed KTO to maximize the utility of LLM generations directly rather than maximizing the log-likelihood of preferences inspired from prospect theory (Tversky and Kahneman, 1992). This approach eliminates the need for two preferences for the same input, as it focuses on discerning whether a preference is desirable or undesirable.

### v3

@@ 2.1 Reinforcement Learning from Human Feedback
The typical RLHF pipeline for LLMs generally consists of three key phases: (1) supervised fine-tuning (SFT), (2) preference sampling and reward model training, and (3) reinforcement learning (RL) optimization.
In the SFT stage, RLHF typically starts with a pre-trained language model, which is further refined using supervised fine-tuning on a curated dataset of high-quality human-generated responses(labels). This supervised fine-tuning stage produces an initial model, denoted as $\pi^{SFT}$, which has improved but still not fully aligned with human preferences. This model serves as the baseline for the subsequent RLHF stages. The preference learning stage involves collecting human feedback in the form of preference data. Given pairs of responses generated by $\pi^{SFT}$, human evaluators express a preference for one response over the other. These preferences serve as the foundational feedback for reward model training. The reward model’s objective is to quantify these preferences by assigning a numerical score to each output, effectively converting human feedback into a scalar reward.
One widely used formula in this stage is the Bradley-Terry Bradley and Terry (1952b) (BT), which proposes a mechanism for modeling pairwise comparisons and assigning concrete reward values to the outputs. In the context of RLHF, BT is used to predict human preferences between pairs of responses. The resulting reward function is then used to guide the concrete model’s decision-making by providing feedback on how well each generated response aligns with human preferences.The BT formula proposes that the human preference distribution $p^{*}$ for a given response pair can be expressed as the following formula:
@@ 2.2 Direct Preference Optimization
In RLHF, the process is relatively intricate, involving the training of a reward model and the iterative sampling from the language model’s policy during the training loop. This complexity arises from the need to continuously evaluate and refine the model based on feedback from the reward model, resulting in significant computational demands. Direct Preference Optimization (DPO) offers a streamlined alternative by optimizing the same objective as RLHF but bypasses the explicit need for a separate reward model, thereby reducing the computational costs associated with aligning the LLM.
@@
This summation ensures that $\pi_{r}(y\mid x)$ is a valid probability distribution, but it is impractical to enumerate all possible responses to traverse all possible input x, which requires significant computational resources, especially when handling a large response space. Actually, DPO algorithm mitigates the need for explicit reward model training and sampling from the LM policy. By directly optimizing preferences without the intermediate step of training a reward model, DPO simplifies and reduces the overall computational burden, making it a more efficient approach for language model alignment. To be specific, the objective function in (4) can be rearranged to isolate $r(x,y)$, yielding:
@@
The DPO loss function is derived directly from the reward differences and avoids the need for a separate reward model. By using this approach, the policy $\pi_{\theta}$ can be analytically optimized based on the reward function differences, simplifying the optimization process and eliminating the need for an intermediate reward model. This makes DPO computationally more efficient, as it directly leverages preference data to guide the model’s learning without the added complexity of reward model training.
@@ 3 Research Questions and Variants
Figure 2: RQ0 Why DPO? The figure shows the pipline of RLHF and DPO. DPO derives a closed-form solution for the optimal policy under the RLHF objective, which allows it to reparameterize the reward function in terms of the policy itself, thereby converting RLHF’s multi-stage process of explicit reward modeling and reinforcement learning into a single-stage direct policy optimization on preference data. The symbols used in this figure are consistent across all subsequent figures in this section.
@@
PPO Ouyang et al. (2022a) $r_{\phi}(x,y)-\beta\mathbb{D}_{\mathrm{KL}}\left[\pi_{\theta}(y\mid x)\|\pi_{\text{ref }}(y\mid x)\right]$
DPO Rafailov et al. (2023) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
IPO Azar et al. (2023) $\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\frac{1}{2}\right)^{2}$
VPO Cen et al. (2025) $\{-&\sum_{s=1}^{t}\log\sigma\left(\beta\log\frac{\pi\left(y_{+}^{(s)}\mid x^{(s)}\right)}{\pi_{\text{ref }}\left(y_{+}^{(s)}\mid x^{(s)}\right)}-\beta\log\frac{\pi\left(y_{-}^{(s)}\mid x^{(s)}\right)}{\pi_{\text{ref }}\left(y_{-}^{(s)}\mid x^{(s)}\right)}\right)\\ &\left.+\alpha\beta\underset{x\sim\rho,y\sim\pi_{\text{cal }}(\cdot\mid x)}{\mathbb{E}}\left[\log\pi(y\mid x)-\log\pi_{\text{ref }}(y\mid x)\right]\right\}\end{aligned}$
@@
RRHF Yuan et al. (2023) $\sum_{r_{i}<r_{j}}\max(0,p_{i}-p_{j})-\sum_{t}\log P_{\pi}(y_{i^{\prime},t}|x,y_{i^{\prime},<t})$
@@
SPIN Chen et al. (2024a) $\ell\left(\lambda\log\frac{p_{\theta}(y|x)}{p_{\theta_{t}}(y|x)}-\lambda\log\frac{p_{\theta}(y^{\prime}|x)}{p_{\theta_{t}}(y^{\prime}|x)}\right)$
@@
Step-DPO Lai et al. (2024) $-\log\sigma(\beta\log\frac{\pi_{\theta}(s_{w}|x,s_{1\sim k-1})}{\pi_{ref}(s_{w}|x,s_{1\sim k-1})}-\beta\log\frac{\pi_{\theta}(s_{l}|x,s_{1\sim k-1})}{\pi_{ref}(s_{l}|x,s_{1\sim k-1})})$
T-DPO Zeng et al. (2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
KTO Ethayarajh et al. (2024) $-\lambda_{w}\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-z_{\text{ref}}\right)-\lambda_{l}\sigma\left(z_{\text{ref}}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right),\,$
@@
cDPO Lin et al. (2024d) $\ell_{c\mathrm{DPO}}=-\sum_{i=1}^{M}\log\sigma(\phi(x_{i},y_{i}^{p})-\phi_{s}(x_{i},y_{i}^{n},s_{i}^{n}))$
$\text{where}\,\,\phi_{s}\left(x_{i},y_{i}^{n},s_{i}^{n}\right)=\gamma\sum_{t=1}^{T}(1-s_{t}^{n})\cdot\log\frac{\pi_{\theta}\left(y_{t}^{n}\left|x_{i},y_{<t}^{n}\right.\right)}{\pi_{\mathrm{ref}}\left(y_{t}^{n}\left|x_{i},y_{<t}^{n}\right.\right)}$
$\text{and}\,\,s_{t}^{n}=\frac{P_{p}(y_{t}^{n}|x_{i},y_{<t}^{n})^{1+\beta}}{P_{n}(y_{t}^{n}|x_{i},y_{<t}^{n})^{\beta}\cdot Z}$
M-DPO Xiong et al. (2025) $\log\sigma\left(\eta\sum_{h=1}^{H}\left[\log\frac{\pi_{\theta,h}\left(a_{h}^{l}\mid s_{h}^{l}\right)}{\pi_{\mathrm{ref},h}\left(a_{h}^{l}\mid s_{h}^{l}\right)}-\log\frac{\pi_{\theta,h}\left(a_{h}^{w}\mid s_{h}^{w}\right)}{\pi_{\mathrm{ref},h}\left(a_{h}^{w}\mid s_{h}^{w}\right)}\right]\right)$
@@
$\beta$-DPO Wu et al. (2024a) $-\log\sigma\left(\beta_{i}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{i}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
CPO Xu et al. (2024b) $-\log p_{\theta}(y_{w}|x)-\log\sigma\left(\beta\log\pi_{\theta}(y_{w}|x)-\beta\log\pi_{\theta}(y_{l}|x)\right)$
ORPO Hong et al. (2024) $-\log p_{\theta}(y_{w}|x)-\lambda\log\sigma\left(\log\frac{p_{\theta}(y_{w}|x)}{1-p_{\theta}(y_{w}|x)}-\log\frac{p_{\theta}(y_{l}|x)}{1-p_{\theta}(y_{l}|x)}\right),\,$
@@
XPO Xie et al. (2025) $\alpha\sum_{i=1}^{t}\log\pi\left(\widetilde{\tau}^{(i)}\right)-\sum_{\left(\tau_{+},\tau_{-}\right)\in\mathcal{D}_{\text{pref }}^{(t)}}\log\left[\sigma\left(\beta\log\frac{\pi\left(\tau_{+}\right)}{\pi_{\text{ref }}\left(\tau_{+}\right)}-\beta\log\frac{\pi\left(\tau_{-}\right)}{\pi_{\text{ref }}\left(\tau_{-}\right)}\right)\right]$
@@
OPtune Chen et al. (2024b) $-R(x,y_{w},y_{l})\cdot\log\sigma\left(\beta_{1}\log\frac{\pi_{t+1}(y_{w}|x)}{\pi_{t}(y_{w}|x)}-\beta_{1}\log\frac{\pi_{t+1}(y_{l}|x)}{\pi_{t}(y_{l}|x)}\right)$
@@
IRPO  Pang et al. (2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(c_{w},y_{w}|x_{i})}{\pi_{t}(c_{w},y_{w}|x_{i})}-\beta\log\frac{\pi_{\theta}(c_{l},y_{l}|x_{i})}{\pi_{t}(c_{l},y_{l}|x_{i})}\right)-\alpha\frac{\log\pi_{\theta}(c_{w},y_{w}|x_{i})}{|c_{w}|+|y_{w}|}$
@@
R-DPO Park et al. (2024) $-\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}-\left(\alpha|y_{w}|-\alpha|y_{l}|\right)\right)$
LD-DPO Liu et al. (2024d) $-\log\sigma\left(\beta_{i}\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta_{i}\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)$
@@
SimPO Meng et al. (2024) $-\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}|x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}|x)-\gamma\right)$
POWER-DL Rashidinejad and Tian (2025) $\mathbb{E}_{\mathcal{D}}\left[l_{i}^{t}L_{\mathrm{POWER}}(x,y_{i}^{+},y_{i}^{-})+(1-l_{i}^{t})L_{\mathrm{POWER}}(x,y_{i}^{-},y_{i}^{+})\right]$
$\text{where}\,\,L_{\mathsf{POWER}}(x,y^{+},y^{-})&:=\log\Big{(}\sigma\Big{(}\beta\left[w(y^{+})\log\pi_{\theta}(y^{+}|x)-w(y^{-})\log\pi_{\theta}(y^{-}|x)+w(y^{+})-w(y^{-})\right]\Big{)}\\ &\qquad+\eta\beta w(y^{+})\log\pi(y^{+}|x)\Big{)}\end{aligned}$
@@
SPO Lou et al. (2024a) $-\log\sigma\left(\xi_{2}\phi_{2}(x,y_{w},y_{l})-\xi_{1}\phi_{1}(x,y_{w},y_{l})\right)$
@@
Table 1: Part variants of DPO objectives discussed in RQs. Note that a method may be discussed across multiple RQs. The table is inspired from  Meng et al. (2024).
@@
RQ0: why DPO? Large Language Models have recently developed rapidly with the supervised fine-tuning (SFT) training Chung et al. (2022). However, due to the next token prediction misaligned with the objective "follow the user’s instructions helpfully and safely" Brown et al. (2020); Fedus et al. (2022), SFT models often exhibit unintended behaviors. For example, they might generate toxic, harmful, hallucinatory, and biased responses Chung et al. (2022). Furthermore, SFT solely instructs models to learn from correct responses while neglecting to address incorrect ones, thus lacking the consideration of human values or preferences Zhao et al. (2023); Ren and Sutherland (2025).
@@
A potential solution to this issue is Reinforcement Learning from Human Feedback (RLHF), using human preferences as a reward signal to fine-tune models by reinforcement learning Chung et al. (2022); Bai et al. (2022b); Lee et al. (2023), such as PPO Schulman et al. (2017b) and actor-critic Mnih et al. (2016); Haarnoja et al. (2018). We follow the mainstream work Ziegler et al. (2019); Stiennon et al. (2020) and focus on PPO Schulman et al. (2017b) in this paper.
@@
Nevertheless, RLHF requires loading four models (policy, reference, reward, and critic) and involves training sensitive hyperparameters along with substantial and costly human oversight to develop a well-trained reward model. Furthermore, some research has identified several challenges associated with reward modeling. For instance, reward misspecification  Pan et al. (2022) occurs when a scalar score from the reward model fails to comprehensively represent human preferences. Additionally, misgeneralization  Tien et al. (2023) arises when reward models compute rewards using unexpected or contingent features of the environment  Michaud et al. (2020), leading to causal confusion and poor out-of-distribution generalization. Another issue is reward hacking  Casper et al. (2023), for example, without regularization to penalize the KL divergence between a base model and the fine-tuned model, large language models (LLMs) undergoing RL often learn to output nonsensical text  Stiennon et al. (2020).
@@
A powerful alternative that does not require an explicit reward model or RL to RLHF is Direct Preference Optimization (DPO) Rafailov et al. (2023). DPO derived the closed-form solution of the PPO optimization objective, revealing the relationship between the reward and the optimal policy model. It then reparameterizes the reward modeling loss with the policy, thereby calculating implicit rewards and optimizing the policy model $\pi_{\theta}$ over the preference data as depicted in EQ.8.
@@
Compared to RLHF, DPO is empirically stable, performant, and computationally lightweight Rafailov et al. (2023); Ethayarajh et al. (2024). However, recent studies indicate that, despite avoiding explicit reward modeling, DPO faces challenges similar to those encountered in RLHF. For example, the offline preferences can lead DPO to overfit the training data distribution, thereby limiting its generalization capabilities Azar et al. (2023). DPO is sensitive to the distribution shift between the base model outputs and preference data. Besides, DPO is prone to generating a biased policy that favors out-of-distribution responses, leading to unpredictable behaviors, etc.  Xu et al. (2024a); Saeidi et al. (2024).
@@
RQ1: Effect of Implicit Reward Modeling. We first discuss the generalization ability of the implicit reward modeling in DPO Wang et al. (2025c). DPO avoids training an explicit reward model by establishing a mapping from reward functions to optimal policies. This enables us to transform a loss function over reward functions into a loss function over policies. $r_{\theta}(x,y)=\beta\log\frac{\pi_{\theta}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}$ is the reward implicitly defined by the policy model $\pi_{\theta}$ and reference model $\pi_{\text{ref}}$.
Figure 3: RQ1 Effect of Implicit Reward Modeling. This figure shows that while the DPO policy distribution fits the in-distribution training data well, its implicit reward mechanism leads to poor generalization. Consequently, its performance degrades significantly on out-of-distribution data, revealing a key limitation under distribution shifts compared to explicit reward models.
@@
IPO Azar et al. (2023) points out the overfitting reward problem in DPO and proposes a novel loss to this issue.  Li et al. (2024a) conducted a theoretic and empirical analysis of the errors inherent in policy optimization methods when learning from user preferences for alignment in reinforcement learning including PPO, DPO, and IPO. Their experimental results underscore the critical importance of optimizing policies using out-of-distribution preference data. Furthermore, the study demonstrates the effectiveness of employing an explicit reward model to enhance policy performance.
@@
Yan et al. (2025) identified the "3D-Properties" in DPO’s implicit reward modeling: Drastic Drop in rejected response likelihood, Degradation into response suppression, and Dispersion effect on unseen responses. Theoretical analysis reveals that gradient imbalance arises from the inverse proportionality between gradient magnitudes of chosen $\pi^{+}$ and rejected $\pi^{-}$ responses, causing gradient explosion for $\pi^{+}$ (as $\pi^{-}\to 0$ ) and vanishing gradients for $\pi^{+}$ Experiments show these issues can be mitigated via adaptive scaling of $\beta$ or integration with SFT loss, though DPO’s sensitivity to preference data distribution remains a fundamental limitation.
Cen et al. (2025) proposed Value-Incentivized Preference Optimization (VPO), whose technical core is the addition of a regularization term to the DPO loss. This term measures the log-probability gain of the current policy over the original reference policy, and a simple sign switch is used to maximize or minimize its expected value. For online scenarios, this mechanism drives exploration by penalizing overconfidence in known behaviors; for offline scenarios, it suppresses overfitting by rewarding consistency with a trusted data distribution.
@@
Figure 4: RQ2 Effect of Different Feedback. The figure contrasts three parallel feedback granularities: Binary (absolute judgment), Pair-wise (relative comparison), and List-wise (multi-item ranking).
Dong et al. (2023) proposed a reward ranked fine-tuning method to explore the list-wise feedback. The core idea of RAFT is that the model iteratively learns from the induced best-of-K policy  Nakano et al. (2022); Cobbe et al. (2021), which samples K responses and selects the one with the highest reward as the final output. Then the model is fine-tuned on the optimal responses.
Rank Responses to align Human Feedback (RRHF)  Yuan et al. (2023) fully exploited rank from human annotators or reward models by combining a modified rank loss with SFT loss. To avoid explicit reward model, they take length-normalized conditional log probability of responses under policy model $\pi_{\theta}$ as reward score. The core idea is to let the policy model $\pi_{\theta}$ give larger probabilities for better responses and give smaller probabilities for worse responses.
@@
Besides,  Ethayarajh et al. (2024) proposed KTO to maximize the utility of LLM generations directly rather than maximizing the log-likelihood of preferences inspired from prospect theory Tversky and Kahneman (1992). This approach eliminates the need for two preferences for the same input, as it focuses on discerning whether a preference is desirable or undesirable.
@@
Figure 5: Step-Level and Token-Level Feedback Methods. The upper part shows Step-DPO, which provides step-level supervision by focusing the optimization unit on the first erroneous reasoning step.The lower part depicts TDPO, which performs finer-grained policy optimization by applying a per-token KL divergence constraint.
Lai et al. (2024) proposed Step-DPO to address the limited effectiveness of DPO on long-chain reasoning tasks. They argued that DPO’s holistic comparison of entire answers lacks the fine-grained supervision needed to identify specific errors within a long reasoning chain. Step-DPO shifts the fundamental unit of optimization from the entire response to an individual reasoning step. The method precisely targets the first erroneous step in a flawed solution and conditioned on the preceding correct steps, trains the model to prefer a correct, self-generated continuation over the original incorrect one. This approach provides more precise process-level supervision, enabling the model to accurately locate and correct errors.